Solve for the unknown using a calculator. Round any decimal answer to the nearest hundredth. Do not use spaces or capital letters.
log (x-47) = 2.3838

Answer by mathdoc314(58)


what is this 'do not use spaces or capital letters'? of course we have to use spaces orallofourwordsrunintoonebiggloblikethis.
log(x-47) = 2.3838
x-47 = 10^(2.3838)
x = 47 + 10^(2.3838)
x = 47 + 288.9914